# Weighted Point Configurations with Hyperuniformity: An Ecological Example and Models

### v1

@@ Abstract
Random point configurations are said to be in hyperuniform states, if density fluctuations are anomalously suppressed in large-scale. Typical examples are found in Coulomb gas systems in two dimensions especially called log-gases in random matrix theory, in which points are repulsively correlated by long-range potentials. In infertile lands like deserts continuous survival competitions for water and nutrition will cause long-ranged repulsive interactions among plants. We have prepared digital data of spatial configurations of center-of-mass for bushes weighted by bush sizes which we call masses. Data analysis shows that such ecological point configurations do not show hyperuniformity as unmarked point processes, but are in hyperuniform states as marked point processes in which mass distributions are taken into account. We propose the non-equilibrium statistical-mechanics models to generate marked point processes having hyperuniformity, in which iterations of random thinning of points and coalescing of masses transform initial uncorrelated point processes into non-trivial point processes with hyperuniformity. Combination of real data-analysis and computer simulations shows the importance of strong correlations in probability law between spatial point configurations and mass distributions of individual points to realize hyperuniform marked point processes.
@@ 1 Introduction
Structures and distribution of ordered and disordered configurations of points in a space have been important research subjects in physics, where the points represent the locations of atoms or molecules in gases, liquids, glasses, and crystals. If we consider a classical ideal-gas, the spatial configurations of molecules are completely random and the statistical ensembles of such uncorrelated point configurations is identified with the Poisson point process (PPP) studied in mathematics [2, 5]. In mathematical physics and probability theory spatial statistics of (random) configurations of points are simply called a (random) point process in general [2, 5]. Correlated point processes have been also extensively studied in statistical mechanics for Coulomb gas systems, which are especially called the log-gases in one- and two-dimensional spaces [6]. It is known that they are are realized as the eigenvalue distributions of Hermitian and non-Hermitian random matrices [4, 6, 7, 10, 19] and as the zeros of random analytic functions [9, 13].
@@
Recently in condensed matter physics and related material sciences, correlated particle systems are said to be in a hyperuniform state, when density fluctuations are anomalously suppressed in large-scale [33]. Consider a bounded domain with linear size $\ell$ in the $d$-dimensional space, which we call an observation window. In this window we measure the mean number of points $\langle N_{\ell}\rangle$ and its variance $\sigma^{2}_{\ell}$. In a hyperuniform system, as the window size $\ell\to\infty$, $\sigma^{2}_{\ell}$ grows more slowly than $\langle N_{\ell}\rangle$ which is proportional to the window volume. Typical disordered systems (e.g., liquids and structured glasses) have the standard scaling behavior $\sigma_{\ell}\sim\ell^{d}$ as $\ell\to\infty$, and hence they are not hyperuniform. Periodic point configuration (e.g., lattice points of crystals) are obviously hyperuniform systems, since the number fluctuations are concentrated near the window boundary and hence have the surface-area scaling $\sigma^{2}_{\ell}\sim\ell^{d-1}$ as $\ell\to\infty$. We are interested in the hyperuniform point processes, which are not perfect lattice, but are randomly distributed following some probability laws. A variety of examples of disordered or random point processes showing hyperuniformity have been reported in the review paper by Torquato [33]. A typical example on the two-dimensional plane, which has been well studied in mathematical physics, is the complex Ginibre point process (GPP) [7]. This is realized as the statistical ensemble of eigenvalues of non-Hermitian random matrices on the complex plane $\mathbb{C}$, which can be identified with the two-dimensional space $\mathbb{R}^{2}$. There first we consider an $n\times n$ matrix whose entries are all independently and identically distributed (i.i.d.) complex standard Gaussian random variables. We have a nearly uniform distribution of eigenvalues in a disk centered at the origin with radius proportional to $\sqrt{n}$ on $\mathbb{C}$. If we consider the $n\to\infty$ limit, we have a translation-invariant point process on $\mathbb{C}$ [7]. It was proved that the obtained GPP is hyperuniform with $\sigma_{\ell}\sim\ell$ [26]. The origin of such hyperuniformity in random point configuration is long-ranged repulsive interaction between any pair of points. As a matter of fact, GPP is a determinantal point process (DPP) [9], for which we can show that a logarithmic repulsive potential acts between any pair of points [6, 18]. We notice that GPP has been studied as a typical example of strongly correlated point process in two dimensions in non-Hermitian physics [1], spectral theory [34], random matrix theory [3], and so on, and used in many applications (see, for instance, [11, 15, 20, 21]).
@@
Figure 1: A sample of point configuration of center-of-masses of bushes in a desert. The sizes of area are indicated not by real lengths in meters but by numbers of pixels for the digital data; $L_{x}=1627$, $L_{y}=895$. The total number of points is $N=9853$ and the density is $\rho\fallingdotseq 6.77\times 10^{-3}$. The bush-size (mass) distribution is represented by colors of each points; the points with masses $1\leq M(\bm{X}_{j})\leq 49$ are green, the points with $50\leq M(\bm{X}_{j})\leq 99$ are blue, and the points with $M(\bm{X}_{j})\geq 100$ are red. See Section 3 how to produce such digital data.
@@
In the present paper, we study the spatial distributions of bushes in deserts. In infertile lands plants can survive only if they have continued to beet the competitions to obtain sufficient amount of water and nutrition. It is expected that such survival competitions will cause strong and long-ranged repulsive interactions among plants. Figure 1 shows one sample of point process obtained from the real data of bushes distributed in the 2195.1 m $\times$ 1208.25 m area in a desert found in the Talampaya Natural Park in Argentina. There we plotted $N=9853$ points $\bm{X}_{j}\in\Lambda$, $j=1,2,\dots,N$, each of which gives the center-of-mass of a bush. As explained in detail in Section 3, we will define a size for each bush, which we call a mass and write as $M(\bm{X}_{j})$, $j=1,2,\dots,N$ in the present paper. We have obtained several real samples of point configurations of center-of-masses of bushes $\mathcal{P}=\{\bm{X}_{j}\}_{j=1}^{N}$ weighted by masses $\{M(\bm{X}_{j}):\bm{X}_{j}\in\mathcal{P}$} in deserts. Such weighted point configurations are called marked point processes in mathematical literature [2, 5]. The partial information of masses is shown by coloring of points in Fig. 1.
@@
In the present paper, we propose statistical mechanics models in non-equilibrium to generate random point configurations weighted by masses on a plane, which do not have hyperuniformity as unmarked point processes, but have hyperuniformity as marked point processes. In our models, we consider a stochastic processes to iterate random thinning of points [16, 17, 31] and coalescing (aggregation, clustering) of masses [8, 22, 25]. Our numerical simulations show that even if we start from PPP, a sufficiently large number of iterations of these two procedures generate non-trivial marked point processes having the desired properties. Comparisons with the real data of bushes are discussed.
@@ 2.1 Marked and unmarked point processes
$\delta_{\bm{x}}(B)=\begin{cases}1,&\quad\mbox{if $\bm{x}\in B$},\\ 0,&\quad\mbox{otherwise}.\end{cases}$ (2.1)
@@
$\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}}\delta_{\bm{X}}(B),$ (2.2)
since (2.2) gives a total number of points included in $B$ for any subset $B\subset S$; $\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}1$. We call $\Xi$ associated with $\mathcal{P}$ a unmarked point process, or simply a point process. We also consider a marked point process, in which each point $\bm{X}\in\mathcal{P}$ carries a variable $M(\bm{X})$. Such a marked point process is denoted by
$\Pi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}}M(\bm{X})\delta_{\bm{X}}(B),\quad B\subset S.$ (2.3)
In this paper, we call $M(\bm{X})$ a mass of the point $\bm{X}\in\mathcal{P}$. For each subset $B\subset S$, (2.3) gives a total mass of the points included in $B$; $\Pi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}M(\bm{X})$. From a marked point process $\Pi$, we can obtain a unmarked point process by deleting the information of masses of points. We denote the unmarked point process obtained from (2.3) as $\Xi_{\Pi}$ in this paper.
@@
Remark 1   If we consider a pair of $\bm{X}_{j}\in S$ and $M(\bm{X}_{j})\in\mathbb{R}$ as a point $\widehat{\bm{X}}_{j}$ in the direct product space $S\times\mathbb{R}$; $(\bm{X}_{j},M(\bm{X}_{j}))\Leftrightarrow\widehat{\bm{X}}_{j}$, then (2.3) can be regarded as a usual point process in $S\times\mathbb{R}$ and written as $\Pi=\sum_{j}\delta_{\widehat{\bm{X}}_{j}}$ [2]. In the present paper, however, we call the point process which can be written in the form (2.3) a marked point process and the point process in the simpler form (2.2) as an unmarked point process in order to clearly distinguish these two types of point processes.
@@ 2.2 Three classes of hyperuniformity
$\mathrm{Var}[\Xi(\Lambda_{\ell})]=\langle(\Xi(\Lambda_{\ell})-\langle\Xi(\Lambda_{\ell})\rangle)^{2}\rangle$ (2.4)
@@
$\mathrm{Var}[\Pi(\Lambda_{\ell})]=\langle(\Pi(\Lambda_{\ell})-\langle\Pi(\Lambda_{\ell})\rangle)^{2}\rangle.$ (2.5)
@@
$R_{\ell}^{\mathrm{point}}:=\frac{\mathrm{Var}[\Xi(\Lambda_{\ell})]}{\langle\Xi(\Lambda_{\ell})\rangle},\quad R_{\ell}^{\mathrm{mass}}:=\frac{\mathrm{Var}[\Pi(\Lambda_{\ell})]}{\langle\Pi(\Lambda_{\ell})\rangle}.$ (2.6)
@@
$R_{\ell}^{\sharp}\simeq\ell^{-1}$ as $\ell\to\infty$. Example 1: GPP $\Xi^{\rm GPP}$ [24, 26, 27, 33]
$S=\mathbb{R}^{2}\simeq\mathbb{C},\qquad\mbox{$\sharp=$ point}.$
@@
$S=\mathbb{R}^{2D}\simeq\mathbb{C}^{D},\quad D\in\mathbb{N}\equiv\left\{1,2,...\right\},\qquad\mbox{$\sharp=$ point}.$
@@
$R_{\ell}^{\sharp}\simeq\ell^{-1}\log\ell$ as $\ell\to\infty$. Example 3: sinc (sine) DPP [4, 19, 28, 29, 30],
$S=\mathbb{R},\qquad\mbox{$\sharp=$ point}.$
Example 4: Euclidean family of DPPs [12]
$S=\mathbb{R}^{d},\quad d\in\mathbb{N},\qquad\mbox{$\sharp=$ point}.$
@@ 3.1 Bush distributions in deserts
A rectangular area was cut out from the Google Maps satellite image of a part of desert. In this example, the rectangular area is an image of a part of the desert centered at the point ($29^{\circ}56^{\prime}02.2^{\prime\prime}\text{S},67^{\circ}50^{\prime}13.6^{\prime\prime}\text{W}$) with length of $2195.1$ m in the east-west (EW) direction and length of $1208.25$ m in the north-south (NS) direction. Bushes are dotted with a variety of sizes as shown in Fig.2 (a).
The image was converted to be gray-scaled using the average method in the computer software Image J. Figure 2 (b) is the result obtained from Fig.2 (a). The resolution of this image is $L_{x}=1626$ pixels by $L_{y}=895$ pixels, which means that each individual pixel corresponds to a geographical area of about $1.35$m $\times$ $1.35$m.
The obtained gray-scaled image was then binarized by automatically setting the threshold value and using the Otsu method, also in Image J. Each pixel is binarized as follows:
@@
The number of black pixels forming a cluster (connected component) was counted as a mass of each bush. By averaging of coordinates of the black pixels included in a cluster, the $x$- and $y$-coordinates of the center-of-mass was calculated for each bush. This procedure was done by Matlab image recognition. Figure 2 (c) shows a part of the obtained table which lists out the identification numbers of the bushes, the masses, the x-coordinates, and the y-coordinates of the center-of-masses.
@@
Table 1: Data of samples studied in this paper. The areas of Nos.1–5 are in the Talampaya Natural Park, La Rioja, Argentina, Nos.5 and 6 in the Parc culturel national de Tindouf, Oum El Assel, Algeria, No. 7 at Flying Fox, Northrm Territory, Australia, and No. 8 at Maikona, Kenya. The latitude and longitude are for the center of each area. Real sizes in the east-west (EW), and north-south (NS) directions are given in the metric unit, whose pixels in the digital data are also shown. The total number of bushes in each area is given by $N$. The areas of Algeria 1, 2 are found in a Wadi and hence the aspect ratios $\lambda=\Lambda_{y}/\Lambda_{x}$ are quite deviated from 1.
@@
We will call the sample shown by Fig, 1 simply (Argentina) bush 1 in the following. By the similar procedure explained above, we have prepared other four samples from the same desert in Argentina, two samples from a desert in Algeria, one from Australia, and one from Kenya. The detailed data are listed out in Table 1.

### v2

@@ Abstract
Random point configurations are said to be in hyperuniform states, if density fluctuations are anomalously suppressed in large-scale. Typical examples are found in Coulomb gas systems in two dimensions especially called log-gases in random matrix theory, in which points are repulsively correlated by long-range potentials. In infertile lands like deserts continuous survival competitions for water and nutrition will cause long-ranged repulsive interactions among plants. We have prepared digital data of spatial configurations of center-of-masses for bushes weighted by bush sizes which we call masses. Data analysis shows that such ecological point configurations do not show hyperuniformity as unmarked point processes, but are in hyperuniform states as marked point processes in which mass distributions are taken into account. We propose the non-equilibrium statistical-mechanics models to generate marked point processes having hyperuniformity, in which iterations of random thinning of points and coalescing of masses transform initial uncorrelated point processes into non-trivial point processes with hyperuniformity. Combination of data analysis and computer simulations shows the importance of strong correlations in probability law between spatial point configurations and mass distributions of individual points to realize hyperuniform marked point processes.
@@ 1 Introduction
Structures and distribution of ordered and disordered configurations of points in a space have been important research subjects in physics, where the points represent the locations of atoms or molecules in gases, liquids, glasses, quasicrystals, and crystals. If we consider a classical ideal-gas, the spatial configurations of molecules are completely random and the statistical ensembles of such uncorrelated point configurations is identified with the Poisson point process (PPP) studied in mathematics [1, 2]. In mathematical physics and probability theory spatial statistics of (random) configurations of points are simply called a (random) point process in general [1, 2]. Correlated point processes have been also extensively studied in statistical mechanics for Coulomb gas systems, which are especially called the log-gases in one- and two-dimensional spaces [3]. It is known that they are realized as the eigenvalue distributions of Hermitian and non-Hermitian random matrices [3, 4, 5, 6, 7] and as the zeros of random analytic functions [8, 9].
@@
Recently in condensed matter physics and related material sciences, correlated particle systems are said to be in a hyperuniform state, when density fluctuations are anomalously suppressed in large-scale [10]. Consider a bounded domain with linear size $\ell$ in the $d$-dimensional space, which we call an observation window. In this window we measure the mean number of points $\langle N_{\ell}\rangle$ and its variance $\sigma^{2}_{\ell}$. In a hyperuniform system, as the window size $\ell\to\infty$, $\sigma^{2}_{\ell}$ grows more slowly than $\langle N_{\ell}\rangle$ which is proportional to the window volume. Typical disordered systems (e.g., liquids and structured glasses) have the standard scaling behavior $\sigma_{\ell}\sim\ell^{d}$ as $\ell\to\infty$, and hence they are not hyperuniform. Periodic point configuration (e.g., lattice points of perfect crystals) are obviously hyperuniform systems, since the number fluctuations are concentrated near the window boundary and hence have the surface-area scaling $\sigma^{2}_{\ell}\sim\ell^{d-1}$ as $\ell\to\infty$. We are interested in the hyperuniform point processes, which are not perfect lattice, but are randomly distributed following some probability laws. A variety of examples of disordered or random point processes showing hyperuniformity have been reported in the review paper by Torquato [10]. A typical example on the two-dimensional plane, which has been well studied in mathematical physics, is the complex Ginibre point process (GPP) [4]. This is realized as the statistical ensemble of eigenvalues of non-Hermitian random matrices on the complex plane $\mathbb{C}$, which can be identified with the two-dimensional space $\mathbb{R}^{2}$. There first we consider an $n\times n$ matrix whose entries are all independently and identically distributed (i.i.d.) complex standard Gaussian random variables. We have a nearly uniform distribution of eigenvalues in a disk centered at the origin with radius proportional to $\sqrt{n}$ on $\mathbb{C}$. If we consider the $n\to\infty$ limit, we have a translation-invariant point process on $\mathbb{C}$ [4]. It was proved that the obtained GPP is hyperuniform with $\sigma_{\ell}\sim\ell$ [11]. The origin of such hyperuniformity in a random point configuration is long-ranged repulsive interaction between any pair of points. As a matter of fact, GPP is a determinantal point process (DPP) [8], for which we can show that a logarithmic repulsive potential acts between any pair of points [3, 12]. It should be remarked that DPP is also called fermion point process [13], since the determinantal structure of correlation functions is essentially equivalent with the Slater determinant describing wave function of fermionic particle systems. We notice that GPP has been studied as a typical example of strongly correlated point process in two dimensions in non-Hermitian physics [14], spectral theory [15], random matrix theory [16], and so on, and used in many applications (see, for instance, the references [17, 18, 19, 20]).
@@
Figure 1: A sample of point configuration of center-of-masses of bushes in a desert. The sizes of area are indicated not by real lengths in meters but by numbers of pixels for the digital data; $L_{x}=1627$, $L_{y}=895$. The total number of points is $N=9853$ and the density is $\rho\fallingdotseq 6.77\times 10^{-3}$. The bush-size (mass) distribution is represented by colors of each points; the points with masses $1\leq M(\bm{X}_{j})\leq 49$ are green, the points with $50\leq M(\bm{X}_{j})\leq 99$ are blue, and the points with $M(\bm{X}_{j})\geq 100$ are red. See Section 3 showing how to produce such digital data.
@@
In the present paper, we study the spatial distributions of bushes in deserts. In infertile lands plants can survive only if they have continued to beat the competitions to obtain sufficient amount of water and nutrition. It is expected that such survival competitions will cause strong and long-ranged repulsive interactions among plants. Figure 1 shows one sample of point process obtained from the real data of bushes distributed in the 2195.1 m $\times$ 1208.25 m area $\Lambda$ in a desert found in the Talampaya Natural Park in Argentina. There we plotted $N=9853$ points $\bm{X}_{j}\in\Lambda$, $j=1,2,\dots,N$, each of which gives the center-of-mass of a bush. As explained in detail in Section 3, we will define a size for each bush, which we call a mass and write as $M(\bm{X}_{j})$, $j=1,2,\dots,N$ in the present paper. We have obtained several real samples of point configurations of center-of-masses of bushes $\mathcal{P}=\{\bm{X}_{j}\}_{j=1}^{N}$ weighted by masses $\{M(\bm{X}_{j}):\bm{X}_{j}\in\mathcal{P}$} in deserts. Such weighted point configurations are called marked point processes in mathematical literature [1, 2]. The partial information of masses is shown by coloring of points in Fig. 1.
@@
In the present paper, we propose statistical-mechanics models in non-equilibrium to generate random point configurations weighted by masses on a plane, which do not have hyperuniformity as unmarked point processes, but have hyperuniformity as marked point processes. In our models, we consider a stochastic processes to iterate random thinning of points [21, 22, 23] and coalescing (aggregation, clustering) of masses [24, 25, 26]. Our numerical simulations show that even if we start from PPP, a sufficiently large number of iterations of these two procedures generates non-trivial marked point processes having the desired properties. Comparisons with the real data of bushes are discussed.
@@ 2.1 Marked and unmarked point processes
$\delta_{\bm{x}}(B)=\begin{cases}1,&\quad\mbox{if $\bm{x}\in B$},\\ 0,&\quad\mbox{otherwise}.\end{cases}$ (1)
@@
$\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}}\delta_{\bm{X}}(B),$ (2)
since (2) gives a total number of points included in $B$ for any subset $B\subset S$; $\Xi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}1$. We call $\Xi$ associated with $\mathcal{P}$ a unmarked point process, or simply a point process. We also consider a marked point process, in which each point $\bm{X}\in\mathcal{P}$ carries a variable $M(\bm{X})$. Such a marked point process is denoted by
$\Pi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}}M(\bm{X})\delta_{\bm{X}}(B),\quad B\subset S.$ (3)
In this paper, we call $M(\bm{X})$ a mass of the point $\bm{X}\in\mathcal{P}$. For each subset $B\subset S$, (3) gives a total mass of the points included in $B$; $\Pi(B)=\sum_{\bm{X}:\bm{X}\in\mathcal{P}\cap B}M(\bm{X})$. From a marked point process $\Pi$, we can obtain a unmarked point process by deleting the information of masses of points. We denote the unmarked point process obtained from (3) as $\Xi_{\Pi}$ in this paper.
@@
Remark 1   If we consider a pair of $\bm{X}_{j}\in S$ and $M(\bm{X}_{j})\in\mathbb{R}$ as a point $\widehat{\bm{X}}_{j}$ in the direct product space $S\times\mathbb{R}$; $(\bm{X}_{j},M(\bm{X}_{j}))\Leftrightarrow\widehat{\bm{X}}_{j}$, then (3) can be regarded as a usual point process in $S\times\mathbb{R}$ and written as $\Pi=\sum_{j}\delta_{\widehat{\bm{X}}_{j}}$ [2]. In the present paper, however, we call the point process which can be written in the form (3) a marked point process and the point process in the simpler form (2) as an unmarked point process in order to clearly distinguish these two types of point processes.
@@ 2.2 Three classes of hyperuniformity
$\mathrm{Var}[\Xi(\Lambda_{\ell})]=\langle(\Xi(\Lambda_{\ell})-\langle\Xi(\Lambda_{\ell})\rangle)^{2}\rangle$ (4)
@@
$\mathrm{Var}[\Pi(\Lambda_{\ell})]=\langle(\Pi(\Lambda_{\ell})-\langle\Pi(\Lambda_{\ell})\rangle)^{2}\rangle.$ (5)
@@
$R_{\ell}^{\mathrm{point}}=\frac{\mathrm{Var}[\Xi(\Lambda_{\ell})]}{\langle\Xi(\Lambda_{\ell})\rangle},\quad R_{\ell}^{\mathrm{mass}}=\frac{\mathrm{Var}[\Pi(\Lambda_{\ell})]}{\langle\Pi(\Lambda_{\ell})\rangle}.$ (6)
@@
$R_{\ell}^{\sharp}\simeq\ell^{-1}$ as $\ell\to\infty$. Example 1: GPP $\Xi^{\rm GPP}$ [10, 11, 27, 28],  $S=\mathbb{R}^{2}\simeq\mathbb{C},\,\,\mbox{$\sharp=$ point}$.
@@
$S=\mathbb{R}^{2D}\simeq\mathbb{C}^{D},\,D\in\mathbb{N}\equiv\left\{1,2,...\right\},\,\,\mbox{$\sharp=$ point}.$
@@
$R_{\ell}^{\sharp}\simeq\ell^{-1}\log\ell$ as $\ell\to\infty$. Example 3: sinc (sine) DPP [5, 6, 13, 29, 30],  $S=\mathbb{R},\,\,\mbox{$\sharp=$ point}$.
Example 4: Euclidean family of DPPs [31],  $S=\mathbb{R}^{d},\,d\in\mathbb{N},\,\,\mbox{$\sharp=$ point}$.
@@ 3.1 Bush distributions in deserts
(i)   A rectangular area was cut out from the Google Maps satellite image of a part of desert. In this example, the rectangular area is an image of a part of the desert centered at the point ($29^{\circ}56^{\prime}02.2^{\prime\prime}\text{S},67^{\circ}50^{\prime}13.6^{\prime\prime}\text{W}$) with length of $2195.1$ m in the east-west (EW) direction and length of $1208.25$ m in the north-south (NS) direction. Bushes are dotted with a variety of sizes as shown in Fig.2 (a).
(ii)   The image was converted to be gray-scaled using the average method in the computer software Image J. Figure 2 (b) is the result obtained from Fig.2 (a). The resolution of this image is $L_{x}=1626$ pixels by $L_{y}=895$ pixels, which means that each individual pixel corresponds to a geographical area of about $1.35$m $\times$ $1.35$m.
(iii)   The obtained gray-scaled image was then binarized by automatically setting the threshold value and using the Otsu method, also in Image J. Each pixel is binarized as follows:
@@
(iv)   The number of black pixels forming a cluster (connected component) was counted as a mass of each bush. By averaging of coordinates of the black pixels included in a cluster, the $x$- and $y$-coordinates of the center-of-mass were calculated for each bush. This procedure was done by Matlab image recognition. Figure 2 (c) shows a part of the obtained table which lists out the identification numbers of the bushes, the masses, the x-coordinates, and the y-coordinates of the center-of-masses.
@@
Table 1: Data of samples
@@
We will call the sample shown by Fig, 1 simply (Argentina) bush 1 in the following. By the similar procedure explained above, we have prepared other four samples from the same desert in Argentina, two samples from a desert in Algeria, one from Australia, and one from Kenya. The detailed data are listed out in Table 1: The areas of Nos.1–5 are in the Talampaya Natural Park, La Rioja, Argentina, Nos.5 and 6 in the Parc culturel national de Tindouf, Oum El Assel, Algeria, No. 7 at Flying Fox, Northrm Territory, Australia, and No. 8 at Maikona, Kenya. The latitude and longitude are for the center of each area. Real sizes in the east-west (EW), and north-south (NS) directions are given in the metric unit, whose pixels in the digital data are also shown. The total number of bushes in each area is given by $N$. The areas of Algeria 1, 2 are found in a Wadi and hence the aspect ratios $\lambda=\Lambda_{y}/\Lambda_{x}$ are quite deviated from 1.